OpenStudy (lukecrayonz):

Part 1: Given log a b = x, replace the variables a and b with integers of your choice. (2 points) Part 2: Show your work to solve the logarithm you created using the change of base formula. (4 points) Part 3: Provide the solution. (2 points)

OpenStudy (anonymous):

\[\log_{2}8=x \]\[2^x=8\]\[2^x=2^3\]\[x=3\]

OpenStudy (anonymous):

that is good but it is not the change of base formula. you can write \[\log_5(100)=\frac{\log(100)}{\log(5)}=\frac{2}{\log(5)}\] and then solve with a calculator

OpenStudy (anonymous):

i missed that re. change of base

OpenStudy (anonymous):

Where did you get the 2 from? @satellite73

OpenStudy (anonymous):

\(\log(100)=2\iff 10^2=100\)